OpenStudy (anonymous):

An open cone has a slant height of 5cm and a cuved surface area of 251cm square .Find the radius of its circular base leaving your answer correct to 3 significant figures

OpenStudy (anonymous):

\[CSA = \pi \times r \times l\] where r is the radius and l is the slant height. \[r=251\div (\pi \times 5)\] =15.9876 cm.

OpenStudy (anonymous):

no u r wrong

Directrix (directrix):

The area of the curved (lateral) surface of a cone = πrL where r is the base radius, and L and L is the slant height. 251 = pi*r*5 251 = 3.1415 * r * 5 251= 15.7075 * r r = ----------> @JudeSharp009 Although you said @NeetziD answer is wrong, I think you can get the same result when you divide. Don't forget this: leaving your answer correct to 3 significant figures